Two very large cruise ships each have a mass of 415 000 tonnes. If they were docked side by side, 1.0 m apart, their centres would be 60 m apart. What is the force of gravitational attraction putting them together? (Hint: 1 tonne = 1000 kg)



1
Expert's answer
2021-07-13T13:33:57-0400
"F_g=\dfrac{Gm_1m_2}{r^2},""F_g=\dfrac{6.67\cdot10^{-11}\ \dfrac{N\cdot m^2}{kg^2}\cdot(4.15\cdot10^8\ kg)^2}{(60\ m)^2}=3191\ N."